OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

A. 1 B. 4 C. 5 D. 25

OpenStudy (chihiroasleaf):

use distributive property then simplify the equation

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+2x=2+6-3 5x=5 x=1